OpenStudy (anonymous):

Which of the following expressions is equal to 3x squared +27

OpenStudy (freckles):

you could factor 3x^2+27 both terms in 3x^2+27 have a common factor 3

OpenStudy (anonymous):

Which of the following expressions is equal to \[x^{2}+9\]

OpenStudy (freckles):

it is hard to tell since there are no expressions given for either 3x^2+27 or x^2+9 but 3x^2+27 is definitely not x^2+9 however 3x^2+27 is a multiple of x^2+9

OpenStudy (anonymous):

well the choices are (x+3i) (x-6i) (x+3i) (x+3i) (x-31) squared (x+3i) squared

OpenStudy (freckles):

so you want to factor x^2+9 well if you have x^2+a^2 then you factor this as (x-ai)(x+ai)

OpenStudy (freckles):

I think you are missing a choice

OpenStudy (freckles):

or a choice has been incorrectly written

OpenStudy (anonymous):

(x-3i) squared is the (x-31) squared

OpenStudy (freckles):

x^2+a^2=(x-ai)(x+ai) so x^2+9=( )( ) use the above formula

OpenStudy (anonymous):

What is the result of isolation \[x^{2}\] in the equation below? \[x^{2}\] + \[(y-5)^{2}\] = 30